OpenStudy (anonymous):

argh. what am i doing wrong here??? line integrals... Evaluate the line integral ∫4xy^6ds, where C is the right half of the circle x^2+y^2=25.

OpenStudy (anonymous):

i have x=5cost and y=5sint. ds=sqrt(5cost^2+5sint^2)dt=5dt. plug that in i have \[\int\limits_{-\pi/2}^{\pi/2}4(5cost)(5sint)^65dt\] which equals 62500/7 but webwork says no..........what?